What is the net ionic equation for 2NH4CL(aq)+Ca(NO3)2(aq)= CaCl2(aq)+2NH4NO3(aq)
Answer #1

Balanced molecular equation:

2NH4Cl(aq)+Ca(NO3)2(aq)= CaCl2(aq)+2NH4NO3(aq)

Total ionic equation:

2NH4+(aq)+2Cl-(aq) + Ca2+(aq) + 2NO3-(aq) = Ca2+(aq) + 2Cl-(aq) + 2NH4+(aq)+ 2NO3-(aq)

The ions on the right side of the equation is same as the ions on the left side of the equation. The reaction does not yield any solid precipitate (e.g. BaSO4(s)) or pure liguid (e.g. H2O(l)) or gaseous product (e.g. CO2(g)). All of the ions in the reaction are spectator ions.

Therefore, no net reaction has happened and no net equation can be written.